(3x^2 +1)/3x^2 = 1 + 1 = 2 ????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (3x^2+1)/3x^2= (3x^2/3x^2)+(1/3x^2) = 1+(1/3x^2)≠ 2 = RHS
The correct statement is (3x^2+1)/3x^2= 1+(1/3x^2)

Thanks